OpenStudy (anonymous):

how do you factor 2h^3+ 7h^2k+ 15hk^2

OpenStudy (anonymous):

take h common =h(2h^2+7hk-15k^2) =h(2h^2+10hk-3hk-15k^2) =h(2h(h+5k)-3k(h+5k)) =h(h+5k)(2h-3k)